In a test of the hypothesis H0: μ=10 versus Ha: μ≠10 a sample of =50 observations possessed mean overbar x=10.6 and standard deviation s=2.6 Find and interpret the​ p-value for this test

The​ p-value for this test is __________. (Round to four decimal places as​ needed.)

Interpret the result. Choose the correct answer below.

A. There is sufficient evidence to reject H0 for α > 0.11.

B.There is insufficient evidence to reject H0 for α=0.15.

C.There is sufficient evidence to reject H0 for α < 0.11.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 10

Ha : \mu   \neq 10

\bar x = 10.6

\mu = 10

s = 2.6

n = 50

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (10.6 - 10) / 2.6 / \sqrt 50

= 1.63

Test statistic = 1.63

Degrees of freedom = n - 1 = 50 - 1 = 49

P-value = 0.1095

The​ p-value for this test is 0.1095 .

\alpha = 0.11

P-value < \alpha

Reject the null hypothesis .

C.There is sufficient evidence to reject H 0 for α < 0.11.